Problem 29.62 A 65.0-cm-diameter cyclotron uses a 550 V oscillating potential difference between the dees. Part A What is the maximum kinetic energy of a proton if the magnetic field strength is 0.820 T Express your answer with the appropriate units. Value J Submit My Answers Give Up Part B How many revolutions does the proton make before leaving the cyclotron? Submit My Answers Give Up
Answer #1

here q = 1.6 x 10-19 C m = 1.672 x 10-27 Kg d = 65 cm R = 0.65 / 2 = 0.325 m   

B = 0.820 T delta V = 550 volt

as we know that the maximum K.E.

K.E. = q x delta V

= 1.6 x 10-19 x 550 = 880 x 10-19   J Ans

as we know that the speed of particle

V = qBR / m

= ( 1.6 x 10-19 x 0.820 x 0.325 ) / ( 1.672 x 10-27 )

= 0.2550 x 108 m/s

here V = wR

w = ( 0.2550 x 108 ) / 0.325

= 0.7846 x 108  rad / s Ans